If y = -x + 1
if x = -3 would it be y = 9 + 1 or y = 10?

Answer by ikleyn(52781)

.
If y = -x + 1
if x = -3 would it be y = 9 + 1 or y = 10?
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

If y = -x + 1  and  x = -3,  then

   y = -(-3) + 1 = 3 + 1 = 4.


The answer to your question is "No".